O A reduce first round consumption by $160 billion: reduce AD by $640 billion O B reduce c O C reduce consumption by $40 billion, reduce AD by $160 billion by $120 billion; reduce AD by $480 billion reduce AD by $160 billion
Answer #1

We are given MPC= 0.75

Also government purchases reduce by $40 billion. This means that the income or total output has reduced by $40 billion. Since MPC is 0.75 , this implies that the change in consumption when income reduces by $40 billion is 0.75 of $40 billion or 0.75*40=$30 billion.

So consumption reduces by $30 billion.

GDP multiplier is given by 1/(1-MPC) = 1/MPS= 1/(1-0.75) = 1/0.25 = 4

So the GDP will reduce by GDP multiplier * change in government purchases

Reduction in AD or GDP = 4*40 = $160 billion

so answer is option D) reduce consumption by $30 billion and AD by $160 billion.

(You can comment in case of doubts )